OpenStudy (anonymous):

Use the squeeze theorem to determine: LIM: x --> infinity [3-sin(e^x)]/[ V(X^2 + 2) can be written as [3 minus Sine (exp x)] divided by [ x- squared plus 2]

terenzreignz (terenzreignz):

Greetings, user ^_^ Welcome to OpenStudy. Without further ado, allow me to explain the squeeze theorem (not in detail, though ;) ) \[\Large \lim_{x\rightarrow \infty}\frac{3-\sin(e^x)}{x^2+2}\] First, however, is that^ the correct expression?

OpenStudy (anonymous):

yes, it is the corerct expression

terenzreignz (terenzreignz):

Well then, we consider this function... \[\Large f(\color{red}x) = \frac{3-\sin(e^\color{red}x)}{\color{red}x^2 + 2}\] And look for functions to 'squeeze' it with. Can you think of a function that will always be less than this function, no matter what value x takes?

terenzreignz (terenzreignz):

Hey, @MERTICH stay with me :>

OpenStudy (anonymous):

ok,

OpenStudy (anonymous):

-1<sin(e^x<1 because i raised e to 100, and the sin is 0.999,

OpenStudy (anonymous):

@terenzreignz hope you r still there

terenzreignz (terenzreignz):

I could say the same thing about you... -_- And I can see where you're going with this, so, let's just be a little more relevant... \[\Large f(\color{red}x) = \frac{3\color{blue}{-\sin(e^\color{red}x)}}{\color{red}x^2 + 2}\] What's the lowest value \(\Large -\sin(e^x)\) could possibly take?

OpenStudy (anonymous):

I think its -1, I mean , sin x can always be between 1 and -1. correct me

terenzreignz (terenzreignz):

That is correct ^_^Therefore, f(x) can never be less than ... \[\Large f(\color{red}x) = \frac{3-\sin(e^\color{red}x)}{\color{red}x^2 + 2} \ge \frac{3\color{green}{-1}}{x^2+2}=\frac{2}{x^2+2}\] Right? ^_^

OpenStudy (anonymous):

so in other words as f(x) approaches infinity, it becomes 0. If you can please give a brief desription of what just happened there, and if possible, draw closer to the equation f(x)=h(x)=g(x).

terenzreignz (terenzreignz):

Hold your horses... we haven't set it up yet, we have only ever found a lower bound for your function... what about an upper bound?

OpenStudy (anonymous):

i think instead of 3-1 we put 3+1 on the numerator from -(-1), to make 4/( x^2 + 2) ??

OpenStudy (anonymous):

still waiting for the upper bound, @terenzreignz

terenzreignz (terenzreignz):

You already got it... I thought you could take it from there... Now... \[\Large \frac{2}{x^2+2}\le \frac{3-\sin(e^x)}{x^2+2}\le\frac{4}{x^2+2}\] Can you squeeze it now?

OpenStudy (anonymous):

yes, i think. because the power on the denominator of the upper bound is greater, as x--> infintiy , it become zero, and the same applies to the lower bound equation, thus, when SQUEEZED, LIM X--> INFINITY, f(x) =0. thank you very much, you have given me more than an answer!

terenzreignz (terenzreignz):

No problem ^_^ Promise me you'll practice, though :)

OpenStudy (anonymous):

Yep, i am still new to the squeeze thoerem, And I will get there soon.